# A novel 3D food printing technique: achieving tunable porosity and fracture properties via liquid rope coiling

### v1

@@ 1 Introduction
3-Dimensional food printing (3DFP) is a rapidly advancing field receiving increasing interest from both academia and the food industry. Extrusion-based printing is the most used method in 3DFP, where a food product is based on a digital design, printed layer by layer until the full design is produced [1, 2]. 3DFP provides new food processing opportunities that are unattainable through conventional food processing methods, which facilitates the creation of personalized products that suit consumers’ taste, textural and nutritional needs [3]. This can be especially useful for consumers with specific needs like dysphagia patients, athletes, and children, or for consumers who want to try new organoleptic experiences such as pieces of chocolate breaking apart in unconventional ways [1, 4, 5, 6].
@@
Fluid instabilities that are intrinsic to the printing process are often among the major challenges involved in 3DFP and result in lower printing resolution or structural inconsistencies [7]. These instabilities can occur because of a myriad of different reasons but are often due to inadequate configuration of the printer or inconsistencies in the filament. One such instability stems from a well-studied phenomenon known as liquid rope coiling (LRC), which occurs when the nozzle height is larger than a critical value, leading to the coiling of the falling stream of viscous liquid upon hitting the substrate [8, 9]. In daily life, LRC is typically observed when pouring honey on toast [10], and has been thoroughly investigated in recent decades [11, 12, 13, 14, 15, 16, 17].
@@
To avoid disruptive effects originating from LRC in extrusion-based 3D printing, the nozzle height is usually minimized. Yet, despite its often undesired nature in 3D printing, LRC can be exploited as a tool for structural design. In this manner, it has been used to create striking geometric structures across various materials and scales, from coiled molten glass and ceramics [18, 19], to nanofiber pottery sculptures [20], and microfluidic channels and variable stiffness foams [21, 22]. Most recently, LRC was used to create porous elastomer structures with distinct mechanical properties for actuation in soft robotics [23]. Given the potential of utilizing LRC to produce diverse structures with tunable porosity, we aim to leverage LRC in a 3DFP setup to establish a framework for creating food structures with novel geometries and textural properties in a relatively rapid manner.
@@
In this study, we exploited LRC in 3DFP to create coiled food structures using different food-based inks (carrot paste, pea paste, and cookie dough). Different printing conditions, including deposition height and extrusion flow rate were investigated to identify the optimal coiling conditions, which were used to create coiled structures with varying degrees of porosity. The porosity was shown to be relevant for programming the mechanical properties of the coiled structures. This study shows that LRC-3DFP is an innovative food structuring method that holds promise for the manufacturing of products with unique textural and sensorial properties and prompts further research in this field.
@@ 2.1 Preparation of 3DFP inks
Carrot and pea 3DFP inks were prepared by mixing carrot powder obtained from residue pulp after juice extraction and toasted pea flour from milled wrinkled green peas with demineralized water in ratios of 12.77:87.23 and 50:50, respectively. These powder to water ratios were pre-determined to provide pastes with desirable extrudability for 3DFP. Mixing was performed in a speedmixer (DAC 150.1 FV-K, FlackTek) at 3500 rpm for 2 minutes in 185 mL PP100 cups. Furthermore, a printable cookie dough was prepared based on methods described in [24]; see Supplementary Information (SI)-Materials for more details.
@@
Extrusion printing was performed using the ‘SUPREME’ 3D food printer, which was built in-house and developed by TNO (The Netherlands). This printer consists of a syringe-based extruder powered by a stepper extruder motor (maximum driving force of 1400 N), a temperature controllable syringe holder for 30 mL plastic syringes, a piston plunger, and a force sensor. This setup enabled us to vary the nozzle and syringe temperature ($T_{\text{nozzle}},T_{\text{syringe}}$ [°C]), printing height (H [mm]), printing (travel) speed ($U_{xy}$ [mm/s]), piston speed (U [mm/s]), and the volumetric flow rate (Q [$\text{mm}^{3}/\text{s}$]). The latter two parameters were calculated as described in SI-Printing parameters. LRC was controlled by varying H, $U_{xy}$ and Q while maintaining a constant $T_{\text{nozzle}}$ and $T_{\text{syringe}}$ of 25°C and nozzle diameters of 1 and 1.5 mm.
Circular patterns with a diameter of 6 cm were printed in a single layer to examine the coiling behavior. To study the porosity and fracture behavior, bulk structures consisting of printed straight lines with a length of 10 cm were printed in various numbers of layers, resulting in a height of approximately 1.7 cm. The porosity of said structures was adjusted by manipulating the flow rate at constant nozzle speeds, with increased flow rates leading to a higher porosity [23]. All of the above designs were created in Gcode Creator, a slicing software developed by TNO, and were printed with varying printer settings to generate different coiling patterns and or porosities. The designs were printed on top of fitted sheets of baking paper that were assumed to be smooth and planar, which were fastened to the printer platform with magnets to prevent slippage.
Photos and videos of the LRC-3DFP process were captured with a 12MP, 30fps 1080p cellphone camera. In some experiments, the videos captured a lower number of frames per second, which was accounted for in the later calculations. Additional images were captured with a Sony A7 II DSLR camera. For the single-pass designs, ImageJ was used to measure the coil diameters (2R) and trackwidth, with averages of five measurements taken per sample. For designs that were printed in multiple layers, the aforementioned parameters were measured using calipers. The angular coiling frequency ($\Omega$) describes the time required for the formation of one coil and was established by analyzing video sequences of the coil formations in Blender. By counting the number of frames required for the formation of up to five consecutive coils, ($\Omega$) of a single loop was established. It is important to note that we assumed the coils to be perfectly circular with a radius of R for our calculations, while they often deviated from the perfect circle and formed slightly teardrop-shaped patterns. A zoom in of the printing of the coiled samples and relevant parameters is presented in Figure 1a.
@@ 2.3 Post-processing of printed structures
Pea and cookie structures printed with a 1.5 mm nozzle were either baked in a convection oven (Leventi Bakermat) at 160°C for 6 minutes, or in an air fryer (Philips) at 160°C for 3 minutes. These baking times were required to produce fully baked structures with moisture contents of 2.91 ± 0.16 and 2.65 ± 0.74 % wb for cookie and pea ink, respectively. These baking times were halved for the 1.0 mm nozzle. Prior to baking, the printed structures were stored in closed containers at a refrigerated temperature of 5 ± 1°C to prevent water evaporation. This storage method was maintained for up to one day to minimize the spreading of the inks during baking.
@@ 2.4 Intrinsic ink properties
The density ($\rho$) of the inks was determined by dividing the mass ($m$) of the ink by its known volume ($V$), as expressed by $\rho=\frac{m}{V}$. The water activity ($A_{w}$) of the inks and baked structures was measured with an Aqualab dew point activity meter. In order to establish the material’s moisture content (MC), approximately 1 g of sample was weighed and was left in an oven at 104°C for at least 24h for the moisture to evaporate, after which it was cooled down to room temperature in a desiccator for an additional 20 min. MC was expressed in terms of % on wet base. Both $A_{w}$ and MC measurements were performed in duplicate.
The rheological properties of the inks were analyzed with an Anton Paar Rheometer MCR302 fitted with a serrated 25 mm plate-plate configuration and a gap distance of 1 mm. Approximately 2 mL of ink was used per measurement, and all excess material was removed from the geometry prior to analysis. Steady shear measurements were taken, and the shear stress was recorded for shear rates between $0.0001$ and $1\ \text{s}^{-1}$ and back. $1\ \text{s}^{-1}$ was chosen as the maximum shear rate because the materials tended to fracture and leave the boundaries of the geometry at higher shear rates. The resulting curves were fitted to the Herschel-Bulkley model, a common method to determine the yield stress $\sigma_{\text{yield}}$ (Dinkgreve et al., 2016). Elastic (G’) and viscous (G”) moduli were determined by amplitude sweeps in oscillatory mode at a frequency of 0.5 Hz, $0.001$-$100\%$ strain, and were also used to establish the yield stress. All measurements were conducted at $25^{\circ}\text{C}$, all samples were allowed to rest for 3 minutes before rheological testing, and all analyses were performed in duplicate.
@@ 2.7 Instrumental texture analysis
Textural parameters were extracted from the cutting and compression responses, as illustrated in Figure 1c. Hardness is defined as the critical pressure of the highest peak ($P_{c}$), brittleness or fracturability is defined as the strain at the highest peak ($\varepsilon_{c}$), and the effective Young’s modulus is used as a measure of initial stiffness, given by the gradient between the maximum peak and respective strain of the cutting test, following $E_{\text{eff}}=\delta P_{c}/\delta\varepsilon_{c}$ [27, 28]. As $P_{c}$ could not be obtained from the compression tests, strains in the linear regime of $\varepsilon<0.01$ were taken alongside the corresponding $P_{c}$ for the determination of $E_{\text{eff}}$.
Major and meaningful fracture identification
The mechanical responses in both cutting and compression tests exhibited significant fluctuations, as expected from heterogenous fragile structures, resulting in numerous maxima and minima in the force-strain curve. Therefore, we identified two sets of peak values for analysis: i) Major fracture: These are the peak values, maxima, that surpass $n$ neighboring maxima on each side, identified using the ”scipy.signal.argrelextrema” method in Python, with $n$ consistently set to 5 for all samples. These major peaks represent the fractures with significant contributions to the mechanical response (red points in Figure 1c), the differences between which are referred to as $\Delta$strain. These peaks are expected to contribute significantly to the sensory perception. ii) Meaningful fracture: These maxima are higher than a threshold (weakest fracture force) from the next minima. To determine the meaningful fractures, we first calculated the difference in the force response between each maximum and the subsequent minima (fracture force, $dF$). We then discarded $dF$ values smaller than the lowest meaningful critical fracture force, $dF_{\text{min}}$. This value signifies the lowest possible fracture force, setting a threshold for noise in the mechanical responses. Essentially, $dF_{\text{min}}$ can be estimated by cutting a single filament of the printed and baked sample. For the experiments reported in the results section, $dF_{\text{min}}$ was about 0.6 N (for the baked cookie filament with a diameter thickness (d) of 1 mm) and always smaller than 1 N. Therefore, we estimated $dF_{\text{min}}$ to be 1 N for all samples. We then calculated its rescaled value as $df=dF/F_{\text{min}}$. This rescaled fracture force indicates how many times its magnitude exceeds its value in the weakest possible fracture event. Peaks under 1 N were filtered out in all analyses as they are assumed to play an insignificant role in the textural properties (highlighted red area in Figure 1c) [29].
@@
Figure 1: (a) Graphic depiction of the printing and material parameters that are important during coiling. (b) Sample volume determined by measurement of the square outer boundaries. The dashed lines represent the cuts made with the texture analyzer blade for the cutting tests, and the highlighted cube that was formed as a result of the cutting was used for the compression test. (c) Textural properties identified from cutting tests.
@@ 3.1 Rheology and printability of 3DFP inks
First, we characterize the rheological properties of the inks to gain insight into their printability. One ink property that is of main importance during printing is the yield stress ($\sigma_{\text{yield}}$), as the force exerted by the printer must be able to exceed $\sigma_{\text{yield}}$ so it can flow through the nozzle, i.e., it must be printable. In Figures 2a-c, the stress-strain curves obtained from steady shear sweeps are fitted with the Herschel-Bulkley model for pea, carrot, and cookie samples, respectively. Pea ink has the highest $\sigma_{\text{yield}}$ (1.85 kPa), closely followed by cookie (1.55 kPa) and carrot (0.91 kPa). The values of $G^{\prime}$ and $G^{\prime\prime}$ obtained from amplitude sweep tests are presented as a function of shear stress in Figure 2d and as a function of shear strain in Figure 2e. These results show that the cookie and carrot inks show the highest and lowest $G^{\prime}$ and $G^{\prime\prime}$, respectively. We identify similar results for $\sigma_{\text{yield}}$ based on Figure 2d, reported in Supplementary Table 1. The yield stress of all inks except carrot fall outside the range of 500-1500 Pa reported in literature for printable materials [30, 31]. Yet, these values are highly dependent on the capacity of the printer’s stepper motor. During experiments at high flow rates, the force exerted by the motor on the syringe occasionally went up to 600 N, which may be higher than what most food 3D printers are capable of.
Figure 2: Rheological data of the 3DFP inks. Steady shear sweeps of (a) pea, (b) carrot, and (c) cookie inks fit with the Herschel-Bulkley model. Viscoelastic moduli versus (d) shear stress and (e) shear strain; G’=solid symbols, G”=open symbols.
Extended details on rheology of the inks, including a validation of their liquid-like coiling behavior, stability, and large amplitude oscillatory shear (LAOS) responses, are discussed in the SI.

### v2

@@ Abstract
Keywords: Additive manufacturing, 3D food printing, Liquid rope coiling, Texture, Texture analysis, X-ray computed tomography, Rheology, Food fracture
@@ 1 Introduction
3-Dimensional food printing (3DFP) is a rapidly advancing field, receiving increasing interest from both academia and the food industry. Extrusion-based printing is the most used method in 3DFP, where a food product is based on a digital design, printed layer by layer until the full design is produced [1, 2]. 3DFP provides new food processing opportunities that are unattainable through conventional food processing methods, which facilitates the creation of personalized products that suit consumers’ taste, textural, and nutritional needs [3]. This can be especially useful for consumers with specific needs like dysphagia patients, athletes, and children, or for consumers who want to try new organoleptic experiences such as pieces of chocolate breaking apart in unconventional ways [1, 4, 5, 6].
@@
Fluid instabilities that are intrinsic to the printing process are often among the major challenges involved in 3DFP and result in lower printing resolution or structural inconsistencies [7]. These instabilities can occur for a myriad of reasons, but they are often due to inadequate configuration of the printer or inconsistencies in the filament. One such instability stems from a well-studied phenomenon known as liquid rope coiling (LRC), which occurs when the nozzle height is larger than a critical value, leading to the coiling of the falling stream of viscous liquid upon hitting the substrate [8, 9]. In daily life, LRC is typically observed when pouring honey on toast [10], and has been thoroughly investigated in recent decades [11, 12, 13, 14, 9, 15, 8, 16, 17, 18, 19].
@@
The nozzle height is usually minimized to avoid disruptive effects originating from LRC in extrusion-based 3D printing. Yet, despite its often undesired nature in 3D printing, LRC can be exploited as a tool for structural designing. In this manner, it has been used to create striking geometric structures across various materials and scales such as coiled molten glass and ceramics [20, 21], nanofiber pottery sculptures [22], microfluidic channels [23], and variable stiffness foams [24]. Recently, LRC was used to create elastomer structures with graded porosity and distinct mechanical properties for actuation in soft robotics [25, 26]. Given the potential of utilizing LRC to produce diverse structures with tunable porosity, we aim to leverage LRC in a 3DFP setup to establish a framework for creating food structures with novel geometries and textural properties in a relatively rapid manner.
@@
In this study, we exploited LRC in 3DFP to create coiled food structures using different food-based inks (carrot paste, pea paste, and cookie dough). Different printing conditions, including deposition height and extrusion flow rate, were investigated to identify the optimal coiling conditions, which were used to create coiled structures with varying degrees of porosity. The porosity was shown to be relevant for programming the mechanical properties of the coiled structures. This study shows that LRC-3DFP is an innovative food structuring method that holds promise for the manufacturing of products with unique textural and sensorial properties and prompts further research in this field.
@@ 2.1 Preparation of 3DFP inks
Carrot and pea 3DFP inks were prepared by mixing carrot powder obtained from residue pulp after juice extraction and toasted pea flour from milled wrinkled green peas with demineralized water in ratios of 12.77:87.23 and 50:50, respectively; see the Supplementary Information (SI) A.1 for more details. These powder-to-water ratios were pre-determined to provide pastes with desirable extrudability for 3DFP. Mixing was performed in a speedmixer (DAC 150.1 FV-K, FlackTek) at 3500 rpm for 2 minutes in 185 mL PP100 cups. Furthermore, printable cookie dough was prepared based on methods described in [27], further elaborated upon in SI A.1.
@@
Extrusion printing was performed using the ‘SUPREME’ 3D food printer, which was built in-house and developed by TNO (The Netherlands). This printer consists of a syringe-based extruder powered by a stepper extruder motor (maximum driving force of 1400 N), a temperature-controllable syringe holder for 30 mL plastic syringes, a piston plunger, and a force sensor. This setup enabled us to vary the nozzle and syringe temperature ($T_{\text{nozzle}},T_{\text{syringe}}$ [°C]), printing height (H [mm]), printing (travel) speed ($U_{xy}$ [mm/s]), piston speed (U [mm/s]), and the volumetric flow rate (Q [$\text{mm}^{3}/\text{s}$]). The latter two parameters were calculated as described in SI A.2. LRC was controlled by varying H, $U_{xy}$ and Q while maintaining a constant $T_{\text{nozzle}}$ and $T_{\text{syringe}}$ of 25°C and nozzle diameters of 1 and 1.5 mm.
Circular patterns with a diameter of 6 cm were printed in a single layer to examine the coiling behavior (Figure 1a). To study the porosity and fracture behavior, bulk structures consisting of printed straight lines with a length of 10 cm were printed in various numbers of layers, resulting in a height of approximately 1.7 cm. The porosity of said structures was adjusted by manipulating the flow rate at constant nozzle speeds, with increased flow rates leading to a higher porosity ([25]; SI A.2). All of the above designs were created in Gcode Creator, a slicing software developed by TNO, and were printed with varying printer settings to generate different coiling patterns and or porosities. The designs were printed on top of fitted sheets of baking paper that were assumed to be smooth and planar, which were fastened to the printer platform with magnets to prevent slippage.
Photos and videos of the LRC-3DFP process were captured with a 12MP, 30fps 1080p cellphone camera. In some experiments, the videos captured a lower number of frames per second, which was accounted for in the later calculations. Additional images were captured with a Sony A7 II DSLR camera. For the single-pass designs, ImageJ was used to measure the coil diameters (2R) and track width, with averages of five measurements taken per sample. For designs that were printed in multiple layers, the aforementioned parameters were measured using calipers. The angular coiling frequency ($\Omega$) describes the time required for the formation of one coil and was established by analyzing video sequences of the coil formations in Blender (video editor). By counting the number of frames required for the formation of up to five consecutive coils, ($\Omega$) of a single loop was established. It is important to note that we assumed the coils to be perfectly circular with a radius of R for our calculations, while they often deviated from the perfect circle and formed slightly teardrop-shaped patterns. A zoomed-in of the printing of the coiled samples and relevant parameters is presented in Figure 1a.
@@ 2.3 Post-processing of printed structures
Pea and cookie structures printed with a 1.5 mm nozzle were either baked in a convection oven (Leventi Bakermat) at 160°C for 6 minutes, or in an air fryer (Philips) at 160°C for 3 minutes. These baking times were required to produce fully baked structures with moisture contents of 2.91 ± 0.16 and 2.65 ± 0.74 % wet basis (wb) for cookie and pea ink, respectively. These baking times were halved for the 1.0 mm nozzle. Before baking, the printed structures were stored in closed containers at a refrigerated temperature of 5 $\pm$ 1°C to prevent water evaporation. This storage method was maintained for up to one day to minimize the spreading of the inks during baking.
@@ 2.4 Intrinsic ink properties
The density ($\rho$) of the inks was determined by dividing the mass ($m$) of the ink by its known volume ($V$), as expressed by $\rho=\frac{m}{V}$. The water activity ($A_{w}$) of the inks and baked structures was measured with an Aqualab dew point activity meter. In order to establish the material’s moisture content (MC), approximately 1 g of the sample was left in an oven at 104°C for at least 24h for the moisture to evaporate, after which it was cooled down to room temperature in a desiccator for an additional 20 min. MC was expressed in terms of % on wet base. Both $A_{w}$ and MC measurements were performed in duplicate.
The rheological properties of the inks were analyzed with an Anton Paar Rheometer MCR302 fitted with a serrated 25 mm plate-plate configuration and a gap distance of 1 mm. Approximately 2 mL of ink was used per measurement, and all excess material was removed from the geometry prior to analysis. Steady shear measurements were taken, and the shear stress was recorded for shear rates between $0.0001$ and $1\ \text{s}^{-1}$ and back. $1\ \text{s}^{-1}$ was chosen as the maximum shear rate because the materials tended to fracture and leave the boundaries of the geometry at higher shear rates. The resulting curves were fitted to the Herschel-Bulkley model, a common method to determine the yield stress $\sigma_{\text{yield}}$ [28]. Elastic (G’) and viscous (G”) moduli were determined by amplitude sweeps in oscillatory mode at a frequency of 0.5 Hz, $0.001$-$100\%$ strain, and were also used to establish the yield stress. All measurements were conducted at $25^{\circ}\text{C}$, all samples were allowed to rest for 3 minutes before rheological testing, and all analyses were performed in duplicate.
@@ 2.7 Instrumental texture analysis
Textural parameters were extracted from the cutting and compression responses, as illustrated in Figure 1c. Hardness is defined as the critical pressure of the highest peak ($P_{c}$), brittleness or fracturability is defined as the strain at the highest peak ($\varepsilon_{c}$), and the effective Young’s modulus is used as a measure of initial stiffness, given by the gradient between the pressure and respective strain of the cutting test, following $E_{\text{eff}}=\delta P/\delta\varepsilon\approx P_{c}/\varepsilon_{c}$ [31, 32]. As $P_{c}$ could not be obtained from the compression tests, strains in the linear regime of $\varepsilon<0.01$ were taken alongside the corresponding $P$ for the determination of $E_{\text{eff}}$.
2.8 Major and meaningful fracture identification
The mechanical responses in both cutting and compression tests exhibited significant fluctuations, as expected from heterogeneous fragile structures, resulting in numerous maxima and minima in the force-strain curve. We identified two sets of peak values for analysis, (i) major fractures and (ii) meaningful fractures.
Major fractures are the peak values, maxima, that surpass $n$ neighboring data points on each side, identified using the ”scipy.signal.argrelextrema” method in Python, with $n$ consistently set to 5 for all samples. These major peaks represent the fractures with significant contributions to the mechanical response, which are separated by $\Delta\varepsilon$ strain (red points in Figure 1c). Peak forces under 1 N were filtered out in all analyses as they are assumed to play an insignificant role in the textural properties (highlighted red area in Figure 1c) [33].
Meaningful fractures are maxima, which are higher than a threshold (weakest fracture force) from the next minima. To determine the meaningful fractures, we first calculated the difference in the force response between each maximum and the subsequent minima (fracture force, $\Delta F$). We then discarded $\Delta F$ values smaller than the lowest meaningful critical fracture force, $\Delta F_{\text{min}}$. This value signifies the lowest possible fracture force, setting a threshold for noise in the mechanical responses. Essentially, $\Delta F_{\text{min}}$ is equivalent to the highest fracture force for cutting a single filament. For the baked cookie filament with a diameter thickness of 1 mm and 1.5 mm, the highest fracture force ($\Delta F_{\text{min}}$) was 0.6 N and 1 N, respectively (See SI D.2). As typical force detection threshold for human is about 1 N [33], we estimated $\Delta F_{\text{min}}$ to be 1 N for all samples in our analysis. We finally calculated the dimensionless meaningful fracture by $\Delta f=\Delta F/\Delta F_{\text{min}}$.
@@ 2.8 Major and meaningful fracture identification
Figure 1: Overview and methodologies: (a) Graphic depiction of the printing and material parameters that are important during coiling. (b) Sample volume determined by measurement of the square outer boundaries. The dashed lines represent the cutting cross-sections made with the texture analyzer blade for the cutting tests, and the highlighted cubes that were formed as a result of the cutting were used for the compression test. (c) Example of mechanical responses (here upon cutting) and quantification of the textural properties.
@@ 3.1 Rheology and printability of 3DFP inks
First, we characterize the rheological properties of the inks to gain insight into their printability. One ink property that is of main importance during printing is the yield stress ($\sigma_{\text{yield}}$), as the force exerted by the printer must be able to exceed $\sigma_{\text{yield}}$ so it can flow through the nozzle. In Figures 2a-c, the stress-strain curves obtained from steady shear sweeps are fitted with the Herschel-Bulkley model for pea, carrot, and cookie samples, respectively. Pea ink has the highest $\sigma_{\text{yield}}$ (1.85 kPa), closely followed by cookie (1.55 kPa) and carrot (0.91 kPa). The values of $G^{\prime}$ and $G^{\prime\prime}$ obtained from amplitude sweep tests are presented as a function of shear stress in Figure 2d and as a function of shear strain in Figure 2e. These results show that the cookie and carrot inks show the highest and lowest $G^{\prime}$ and $G^{\prime\prime}$, respectively. We identify similar results for $\sigma_{\text{yield}}$ based on Figure 2d, reported in SI Table B.1. The yield stress of all inks except carrot fall outside the range of 500-1500 Pa reported in literature for printable materials [34, 35]. Yet, these values are highly dependent on the capacity of the printer’s stepper motor. During experiments at high flow rates, the force exerted by the motor on the syringe occasionally went up to 600 N, which may be higher than what most food 3D printers are capable of.
Figure 2: Rheological properties of the 3DFP inks: Shear stress response as a function of the applied shear rate upon steady shear sweeps of (a) pea (green), (b) carrot (orange), and (c) cookie (yellow) inks fit with the Herschel-Bulkley model. Viscoelastic moduli versus (d) shear stress and (e) shear strain; G’=solid symbols, G”=open symbols.
Extended details on rheology of the inks, including a validation of their liquid-like coiling behavior, stability, and large amplitude oscillatory shear (LAOS) responses, are discussed in SI B.